OpenStudy (anonymous):

find the area bounded by the curve 4y=x^3 the x-axis and the line x=8

OpenStudy (perl):

did you draw a graph of the region first

OpenStudy (perl):

I graphed it here https://www.desmos.com/calculator/wgpjlgs9c2

OpenStudy (perl):

$$ \text{area between curves f(x) and g(x) is} \\ \int_{a}^{b} f(x) - g(x) $$

OpenStudy (perl):

the top curve is f(x) = x^3/4 bottom curve g(x) = 0 $$ \Large \int_{0}^{8}(\frac{x^3}{4} - 0) ~dx $$

OpenStudy (anonymous):

thank you :)